# LAPD: Langevin-Assisted Bayesian Active Learning for Physical Discovery

### v1

@@ Abstract
Discovering physical laws from data is a fundamental challenge in scientific research, particularly when high-quality data are scarce or costly to obtain. Traditional methods for identifying dynamical systems often struggle with noise sensitivity, inefficiency in data usage, and the inability to quantify uncertainty effectively. To address these challenges, we propose Langevin-Assisted Active Physical Discovery (LAPD), a Bayesian framework that integrates replica-exchange stochastic gradient Langevin Monte Carlo to simultaneously enable efficient system identification and robust uncertainty quantification (UQ). By balancing gradient-driven exploration in coefficient space and generating an ensemble of candidate models during exploitation, LAPD achieves reliable, uncertainty-aware identification with noisy data. In the face of data scarcity, the probabilistic foundation of LAPD further promotes the integration of active learning (AL) via a hybrid uncertainty-space-filling acquisition function. This strategy sequentially selects informative data to reduce data collection costs while maintaining accuracy. We evaluate LAPD on diverse nonlinear systems such as the Lotka-Volterra, Lorenz, Burgers, and Convection-Diffusion equations, demonstrating its robustness with noisy and limited data as well as superior uncertainty calibration compared to existing methods. The AL extension reduces the required measurements by around 60% for the Lotka-Volterra system and by around 40% for Burgers’ equation compared to random data sampling, highlighting its potential for resource-constrained experiments. Our framework establishes a scalable, uncertainty-aware methodology for data-efficient discovery of dynamical systems, with broad applicability to problems where high-fidelity data acquisition is prohibitively expensive.
@@ 1 Introduction
Data-driven discovery of dynamical systems has emerged as a transformative paradigm for extracting governing equations from observational data, enabling insights into complex phenomena in physics (Brunton et al.,, 2016), biology (Zheng et al.,, 2022) and engineering (Kaheman et al.,, 2020). Advances in sparse regression techniques, such as Sparse Identification of Nonlinear Dynamics (SINDy) (Brunton et al.,, 2016; Rudy et al.,, 2016), and Neural Networks (Champion et al.,, 2019; Raissi et al.,, 2019), have demonstrated success in identifying parsimonious models under idealized conditions. Subsequently, several extensions have been developed to address real-world challenges (Messenger and Bortz,, 2021; Zhang and Lin,, 2018). However, real-world applications often face critical challenges: Experimental data are frequently corrupted by noise, scarce because of high acquisition costs, or sampled sparsely in high-dimensional domains. Traditional methods, which rely heavily on deterministic optimization or fixed regularization, struggle to quantify uncertainty in identified models, exhibit sensitivity to noise (Zheng and Lin,, 2024), and demand large datasets for reliable performance, which limits their utility in resource-constrained settings.
Bayesian methods, such as stochastic gradient Langevin dynamics (SGLD) and Hamiltonian Monte Carlo, leverage gradient-driven dynamics to explore posterior distributions of learned coefficients efficiently (Welling and Teh,, 2011; Neal et al.,, 2011). Although these methods improve traditional Markov Chain Monte Carlo (MCMC) techniques, they often face challenges when identifying convex dynamical systems with multi-modal regions, such as local mode trapping or sensitivity to identified parameters (Hirsh et al.,, 2022). To address this, replica exchange stochastic gradient Langevin dynamics (reSGLD) addresses these limitations by combining parallel sampling chains at multiple temperatures (Deng et al.,, 2020; Zheng et al.,, 2024), which enables aggressive exploration of identified coefficient space through high-temperature chains while low-temperature chains refine high-probability regions. This balance is critical for identifying sparse, interpretable dynamical systems from noisy data and providing robust UQ.
In resource-constrained scenarios, AL strategies minimize data acquisition costs by iteratively selecting measurements that maximize information gain. Traditional approaches prioritize either uncertainty reduction (e.g., querying regions with high predictive variance) or space-filling designs (Meng and Karniadakis,, 2020) (e.g., uniform domain coverage). However, purely uncertainty-driven physical discovery risks over-exploiting localized regions, while diversity-focused strategies neglect model confidence. Hybrid acquisition functions, which combine uncertainty and spatial diversity metrics, mitigate this trade-off by selecting points that resolve model ambiguity while ensuring broad coverage of the system’s phase space. For dynamical systems, this might involve targeting states where trajectory predictions diverge across posterior samples or where sparse measurements leave governing equations underdetermined. Bayesian frameworks naturally enable such strategies by quantifying epistemic uncertainty, yet existing methods often lack scalable sampling or adaptive acquisition protocols.
To address such issues, the proposed Langevin-Assisted Active Physical Discovery (LAPD) unifies these advances into a cohesive framework for data-efficient, uncertainty-aware system identification. By balancing exploration and exploitation with reSGLD and a hybrid uncertainty-space-filling acquisition function in AL, LAPD achieves robust posterior estimation even with noisy and limited data. The framework adaptively prioritizes measurements that maximally reduce uncertainty in model coefficients while ensuring diverse sampling of the dynamical landscape—effectively minimizing experimental costs. Our contributions include:
  • a scalable Bayesian framework considering reSGLD for uncertainty-aware system identification;
  • an AL protocol that hybridizes uncertainty and diversity criteria for efficient data sampling;
  • empirical validation across the Lotka-Volterra system, Lorenz system, Burgers’ equation, and convection-diffusion system to validate its accurate UQ and data acquisition.
@@
This work bridges the gap between probabilistic inference and data efficiency, which offers a principled tool for scenarios where high-fidelity data are prohibitively expensive.
@@
The remainder of this paper is organized as follows. Section 2 formulates the problem of data-driven dynamical system discovery. Section 3 introduces the LAPD framework, which details its integration of reSGLD for uncertainty-aware model identification. Section 4 validates the framework through experiments on four non-linear systems: the Lotka-Volterra system, the Lorenz system, the Burgers equation, and the convection-diffusion equation, which demonstrates robustness to noise and data scarcity. In Section 5, we extend LAPD with an AL strategy that leverages its UQ capabilities to optimize data acquisition, with benchmarks on the Lotka-Volterra and Burgers’ systems under limited-data regimes. Finally, Section 6 discusses the broader implications, limitations, and future directions.
@@ 2 Problem Statement
2.1 Physical Discovery
We study a class of differential equations given by:
$\frac{\mathrm{d}}{\mathrm{d}t}\mathbf{u}(t)=\mathbf{f}(\mathbf{u}(t),t),$ (1)
@@ 2.1 Physical Discovery
where the state vector is defined as $\mathbf{u}(t)=[u_{1}(t),u_{2}(t),\cdots,u_{d}(t)]^{\intercal}\in\mathbb{R}^{d}$, and $t$ denotes the corresponding temporal inputs. For simplicity, we exclude spatial inputs from the formulation (1), although such extensions could include partial derivatives with respect to spatial variables. The dynamics of the system is encoded by the function $\mathbf{f}(\cdot)$, which is unknown and usually of non-linear form.
@@
The dynamics of the system $\mathbf{f}(\cdot)$ is learned from the data collected at time $t=t_{n},\ n=1,2,3,\cdots,N$. The size of the data collected is $N$, with the collected data denoted as $\mathbf{U}=[\mathbf{u}(t_{1}),\mathbf{u}(t_{2}),\cdots,\mathbf{u}(t_{N})]^{\intercal}\in\mathbb{R}^{N\times d}$. The learning of $\mathbf{f}(\cdot)$ is accomplished by assuming that the underlying dynamical system can be expressed as a linear combination of functions from a candidate library $\bm{\Theta}(\mathbf{U})=\begin{bmatrix}\theta_{1}(\mathbf{U}),\cdots,\theta_{m}(\mathbf{U})\end{bmatrix}\in\mathbb{R}^{N\times m}$:
$\dot{\mathbf{U}}=\bm{\Theta}(\mathbf{U})\bm{\Xi},$ (2)
@@
where $\dot{\mathbf{U}}\in\mathbb{R}^{N\times d}$ denotes the temporal derivative of the $\mathbf{U}$, $\bm{\Xi}\in\mathbb{R}^{m\times d}$ are coefficients that determine which specific functions from the library $\bm{\Theta}(\mathbf{U})$ contribute to the system dynamics. A typical choice for the candidate functions includes polynomials of the state variables (as well as the partial derivatives of $x$ for PDEs), as they are commonly associated with canonical models of dynamical systems. Fourier-based libraries, incorporating terms such as $sin(\mathbf{x})$ and $cos(\mathbf{x})$, are also frequently employed.
A key assumption to characterize the system is that the coefficient matrix $\mathbf{\Xi}$ is sparse, i.e., the system can be represented by a small number of candidates in the library $\bm{\Theta}$. The standard Sparse Identification of Nonlinear Dynamics (SINDy) (Brunton et al.,, 2016; Rudy et al.,, 2016) uses a sequentially thresholded least-squares (STLSQ) algorithm to perform sparse inference. The STLSQ method serves as a proxy for $l_{0}$ optimization and comes with established convergence guarantees (Zhang and Schaeffer,, 2018). This algorithm combines sparse regression techniques like Ridge and Lasso regression with thresholding to eliminate insignificant terms. To enhance robustness and incorporate UQ, we reformulate the identification of system (1) within a Bayesian framework as follows.
2.2 Bayesian Inference for Data-driven Discovery
@@
$p(\mathbf{\Xi,\Sigma}|\mathbf{U})\propto p(\mathbf{\Xi})p(\mathbf{\Sigma})\prod_{i=1}^{N}p(\mathbf{u}_{i}|\mathbf{\Xi,\Sigma}),$ (3)
@@
where the term $\mathbf{\Sigma}$ encompasses additional parameters of the probabilistic model, including variables such as the standard deviation associated with the random noise term. The computation of the posterior distribution (Equation (3)) typically presents analytical intractability. In such cases, sampling-based approaches, such as Markov Chain Monte Carlo (MCMC) methods, provide a viable solution. The approximated posterior distribution enables model reconstructions and trajectory forecasts given observational data, allowing quantitative uncertainty assessment of model coefficients and predicted trajectories.
Within the Bayesian framework, it necessitates a prior distribution $p(\mathbf{\Xi})$ that promotes sparsity while robustly handling coefficients of varying magnitudes. The Regularized Horseshoe prior (Piironen and Vehtari,, 2017) is particularly well-suited for this task, as it combines strong shrinkage for irrelevant coefficients with controlled regularization for significant ones—avoiding over-shrinkage of true signals common in standard sparsity priors. To adopt this, we first assign each coefficient $\beta_{i}$ a local shrinkage parameter $\tilde{\lambda}_{i}$, which determines how strongly it is shrunk toward zero:
$\beta_{i}|\tilde{\lambda}_{i},\tau,c\sim\mathcal{N}(0,\tilde{\lambda}_{i}^{2}\tau^{2}),\ \ \text{where}\ \ \tilde{\lambda}_{i}=\frac{c\lambda_{i}}{\sqrt{c^{2}+\tau^{2}\lambda_{i}^{2}}}.$

### v2

# BLADE: Bayesian Langevin Active Discovery with Replica Exchange for Identification of Complex Systems

@@ Abstract
Traditional methods for system discovery frequently struggle with efficient data usage and uncertainty quantification. Identifying the governing equations of complex dynamical systems from data presents a significant challenge in scientific discovery, especially when high-quality measurements are scarce and expensive to obtain. To overcome these limitations, we propose Bayesian Langevin Active Discovery with Replica Exchange for Identification of Complex Systems (BLADE), a novel Bayesian framework that combines replica-exchange stochastic gradient Langevin Monte Carlo with active learning. By balancing gradient-driven exploration and exploitation in coefficient space, BLADE provides probabilistic parameter estimation and principled uncertainty quantification. Faced with data scarcity, the probabilistic foundation of BLADE further facilitates the integration of active learning through a hybrid acquisition strategy that combines predictive uncertainty with space-filling design, enabling efficient selection of informative samples. Across benchmark systems, BLADE reduces measurement requirements by roughly 60% for Lotka–Volterra and 40% for Burgers’ equation relative to random sampling, demonstrating substantial data-efficiency gains. These results highlight BLADE as a general uncertainty-aware framework for discovering interpretable dynamical systems, particularly valuable when high-fidelity data acquisition is prohibitively expensive.
@@ 1 Introduction
Identification of governing equations of complex dynamical systems from data is a longstanding goal in science and engineering. These governing equations, expressed as accurate dynamical models, underpin predictive capability and scientific understanding in domains such as turbulence modeling [Pope2000], climate dynamics [drake2014], systems biology [alon2006], and neuroscience [izhikevich2007]. While many systems are known to follow differential equations, deriving their precise forms from first principles can be intractable due to complexity or incomplete knowledge. In such settings, data-driven approaches provide a complementary pathway: they reconstruct the underlying dynamics from measurements, enabling accurate simulation and forecasting.
In recent years, there has been a surge of work on learning nonlinear governing equations from data, which mainly follows two complementary tracks. First, deep-learning-based formulations approximate the dynamics as universal function approximators, where deep neural network frameworks combined with numerical integration schemes provide black-box approaches to approximate nonlinear differential equations from observed data. Qin et al. [Qin2019] and Raissi et al. [Raissi2018] use multi‐step and integral formulations to approximate unknown ordinary differential equations, while Goyal and Benner [Goyal2021] embed a fourth‐order Runge–Kutta constraint to accommodate noisy and sparse observations. Gao et al.[Gao2020] employ a nonintrusive reduced‐order approach to tackle large-scale systems described by partial differential equations, and Wu and Xiu [Wu2020], along with Chen et al. [Chen2022] use evolution operators in modal or nodal spaces to model time‐dependent PDEs. Chen et al. [chen2018neural] introduce continuous-depth models known as neural ordinary differential equations, and deep backward schemes applied by Germain et al. [germain2022approximation] together with high-dimensional approaches from Han et al. [Han2018] address complex PDEs. These deep models emphasize predictive accuracy but often sacrifice interpretability and often uncertainty awareness. In addition to neural formulations, probabilistic nonparametric approaches such as Gaussian process vector fields provide calibrated uncertainty estimates while learning unknown ODE dynamics from sparse observations [Heinonen2018], though interpretability in terms of concise symbolic structure may not be explicit. A second class of methods, built around sparse regression and symbolic discovery, seeks interpretable structure rather than purely predictive power. There are data-driven approaches that learn interpretable models balancing prediction with scientific insight, among which the Sparse Identification of Nonlinear Dynamics (SINDy) [brunton2016discovering, rudy2019data] framework has become a cornerstone. SINDy assumes that governing dynamics can be expressed as a sparse linear combination of candidate functions, and employs sparse regression to identify the active terms. It has emerged as a transformative paradigm for extracting governing equations from observational data, providing insights into complex phenomena in physics [brunton2016discovering], biology [zheng2022data], and engineering [kaheman2020SINDy]. Its success stems from combining interpretability, parsimony, and flexibility in library design.
A critical step in system discovery is the quantification of uncertainty to assess model reliability and compare competing hypotheses. The standard SINDy framework yields a point estimate of the governing equations; in this work, we extend this approach to provide an approximate distribution over the parameters within the selected sparse subspace, thus quantifying the structural and parametric uncertainty. To extend the point estimation of SINDy with a parametric uncertainty guarantee, several extensions have been proposed. Ensemble-based approaches, like Ensemble SINDy [Fasel_2022], improve robustness by aggregating multiple regressions to provide uncertainty estimates, which can also guide active learning strategies. However, these ensemble-based uncertainties are heuristic and do not correspond to formal posterior distributions. In contrast, Bayesian variants such as UQ-SINDy [hirsh2022sparsifying] offer a more principled probabilistic perspective, using sparsity-promoting priors to infer coefficient posteriors. Nevertheless, these methods are often highly sensitive to prior assumptions and can struggle to capture complex, multi-modal posterior landscapes [fung2024]. Consequently, current frameworks either rely on heuristic ensembles lacking statistical grounding or on Bayesian samplers constrained by prior design and sampling inefficiency. Recent extensions further combine sparse discovery with probabilistic or generative modeling to represent stochasticity in governing equations [jacobs2023hypersindy], highlighting growing interest in uncertainty-aware and distributional system identification beyond point estimates.
Well-calibrated UQ through the Bayesian framework also suggests a natural mechanism for experimental design: in resource-constrained scenarios, one can actively exploit uncertainty to guide data acquisition. This motivates the use of active learning (AL), where strategies iteratively select informative measurements to minimize data costs. Traditional AL approaches prioritize either uncertainty reduction (e.g., sampling where predictive variance is high) [MacKay1992, gramacy2009adaptive] or space-filling designs [loeppky2010, yu2010] (e.g., ensuring broad coverage of the state space). Each approach alone has limitations: uncertainty-driven sampling risks over-exploiting local regions, while space-filling designs may waste effort in regions already well understood. For dynamical systems, informativeness is often heterogeneous — states near bifurcations or transients may yield far more insight than equilibrium regions [chen2020]. These challenges motivate hybrid acquisition strategies that combine uncertainty and diversity, ensuring that new samples both resolve ambiguity and enrich dynamical coverage. On the other hand, practical Bayesian discovery frameworks remain limited: scalable posterior sampling is challenging, and existing AL strategies are rarely integrated with full posterior exploration. As a result, most current methods either provide uncertainty without adaptive acquisition [fung2024], or apply AL heuristically without calibrated uncertainty [Fasel_2022]. Thus, an integrated Bayesian-active-learning paradigm capable of both calibrated uncertainty quantification and efficient data acquisition remains elusive.
In the face of the above-mentioned challenges, we propose Bayesian Langevin Active Discovery with Replica Exchange for Identification of Complex Systems (BLADE), a unified framework that integrates posterior sampling with hybrid active learning. BLADE leverages replica-exchange stochastic gradient Langevin dynamics (reSGLD) [deng2020non, zheng2024constrained] to sample from the posterior distribution of coefficients. By combining parallel sampling chains at multiple temperatures, reSGLD enables aggressive exploration of identified coefficient space through high-temperature chains, while low-temperature chains refine high-probability regions. This balance is critical for identifying sparse and interpretable dynamical systems from observational data and providing robust UQ. In contrast to UQ-SINDy, which relies heavily on prior design, BLADE achieves scalable posterior exploration through stochastic gradient sampling, which will be further illustrated in Section 4.1. Building on this probabilistic foundation, BLADE introduces a hybrid acquisition function that unites predictive uncertainty and space-filling design, systematically reducing data requirements while maintaining accuracy. Our contributions are threefold:
Robust Bayesian Uncertainty Quantification for System Discovery.
BLADE leverages reSGLD to efficiently explore high-dimensional coefficient spaces, which overcomes limitations of traditional MCMC methods in exploring multi-modal landscapes and thus enables accurate posterior estimation.
Data Efficient Hybrid Active Learning Strategy.
A combined predictive-variance + maximin-distance criterion guides measurement selection, cutting data requirements by 40–60% across benchmark systems.
Empirical Validation Across Multiple Dynamical Systems.
BLADE achieves state-of-the-art performance in accuracy, uncertainty calibration, and model sparsity across the Lotka–Volterra and Lorenz systems, as well as Burgers’ and convection–diffusion equations. The robustness of the framework is further demonstrated through efficient AL-driven discovery in limited-data regimes.
@@
This work bridges the gap between probabilistic inference and data efficiency, offering a principled tool for scenarios where high-fidelity data are prohibitively expensive.
@@
The remainder of this paper is organized as follows: Section 2 formulates the problem of data-driven dynamical system discovery. Section 3 introduces the BLADE framework, which details its integration of reSGLD for uncertainty-aware model identification and an AL strategy that leverages its UQ capabilities to optimize data acquisition in the face of data scarcity. Section 4 validates the framework through experiments on four non-linear systems: the Lotka-Volterra system, the Lorenz system, the Burgers’ equation, and the convection–diffusion equation, which demonstrates robustness to noise. In Section 5, we extend the experiments with AL, with benchmarks on the Lotka-Volterra and Burgers’ systems under limited-data regimes. Finally, Section 6 discusses the broader implications, limitations, and future directions.
@@ 2 Problem Statement
2.1 Sparse Dynamical System Discovery
We study a class of autonomous differential equations given by:
$\frac{\mathrm{d}}{\mathrm{d}t}\mathbf{u}(t)=\mathbf{f}(\mathbf{u}(t)),$ (1)
@@ 2.1 Sparse Dynamical System Discovery
where the state vector is defined as $\mathbf{u}(t)=[u_{1}(t),u_{2}(t),\cdots,u_{d}(t)]^{\intercal}\in\mathbb{R}^{d}$, and $t$ denotes the corresponding temporal inputs  [^1] For simplicity, we exclude spatial inputs from the formulation (1). It should be noted that such extensions could easily extend to partial derivatives with respect to spatial variables. . The dynamics of the system are encoded by the function $\mathbf{f}(\cdot)$, which is unknown and usually of non-linear form.
@@
For the purpose of identifying the system $\mathbf{f}(\cdot)$, data is collected at a sequence of discrete time instances $\{t_{n}\}_{n=1}^{N}$. The resulting measurements are aggregated into the matrix $\mathbf{U}=[\mathbf{u}(t_{1}),\mathbf{u}(t_{2}),\cdots,\mathbf{u}(t_{N})]^{\intercal}\in\mathbb{R}^{N\times d}$. The learning of $\mathbf{f}(\cdot)$ is accomplished by assuming that the underlying dynamical system can be expressed as a linear combination of functions from a candidate library $\bm{\Theta}(\mathbf{U})=\begin{bmatrix}\theta_{1}(\mathbf{U}),\cdots,\theta_{m}(\mathbf{U})\end{bmatrix}\in\mathbb{R}^{N\times m}$:
$\dot{\mathbf{U}}=\mathbf{\Theta}(\mathbf{U})\mathbf{\Xi},$ (2)
@@
where $\dot{\mathbf{U}}\in\mathbb{R}^{N\times d}$ denotes the temporal derivative of the $\mathbf{U}$, $\mathbf{\Xi}\in\mathbb{R}^{m\times d}$ are coefficients that determine which specific functions from the library $\mathbf{\Theta}(\mathbf{U})$ contribute to the system dynamics. The library of candidate functions is typically constructed from terms known to appear in canonical models of dynamical systems. A common choice is a polynomial basis of the state variables, which, for PDEs, is augmented with partial derivatives. Fourier-based libraries incorporating trigonometric terms, such as $\sin(\mathbf{x})$ and $\cos(\mathbf{x})$, are also frequently employed, especially for systems with periodic behavior.
One key assumption to characterize the system is that the coefficient matrix $\mathbf{\Xi}$ is sparse, i.e., the system can be represented by a small number of candidates in the library $\mathbf{\Theta}$. This is physically justified, as the canonical governing equations describing most natural phenomena are themselves sparse, comprising only a few essential terms. The standard SINDy framework [brunton2016discovering] performs sparse system identification using regularized regression, most commonly via Lasso or sequentially thresholded least squares (STLSQ). In this formulation, the discovery problem is decomposed into $d$ regression tasks, one for each state derivative. This idea was later extended to PDE discovery (PDE-FIND), where sparse coefficients are identified from a candidate library of spatial and nonlinear terms using sequential thresholding with ridge regression [rudy2019data]. In practice, the sequential thresholding step is used to promote parsimony by iteratively removing coefficients with small magnitude and refitting the remaining active bases. This procedure is effective for recovering sparse governing equations and has well-studied convergence behavior in related settings [zhang2019convergence]. However, it is fundamentally a deterministic sparsification mechanism and, by itself, does not provide posterior uncertainty over the inferred coefficients. To incorporate uncertainty quantification, we build on this sparse-identification framework and introduce a Bayesian treatment of the coefficient inference problem. In our approach, posterior sampling is used to characterize uncertainty in the coefficients, while sequential thresholding is retained as a practical support-selection step for model parsimony. To enable rigorous uncertainty quantification and robust inference, we reformulate the identification of system (1) within a Bayesian framework as follows.
2.2 Probabilistic Framework for System Discovery
@@
$p(\mathbf{\Xi}|\mathbf{U})\propto p(\mathbf{\Xi})\prod_{i=1}^{N}p(\mathbf{u}(t_{i})|\mathbf{\Xi}),$ (3)
@@
where $p(\mathbf{u}(t_{i})|\mathbf{\Xi})$ denotes the likelihood for a single observation. The computation of the posterior distribution $p(\mathbf{\Xi}|\mathbf{U})$ typically presents analytical intractability. In such cases, sampling-based approaches, such as Markov Chain Monte Carlo (MCMC) methods, provide a viable solution. The approximated posterior distribution enables model reconstructions and trajectory forecasts given observational data, allowing quantitative uncertainty assessment of model coefficients and predicted trajectories.
Within the Bayesian framework, it necessitates a prior distribution $p(\mathbf{\Xi})$ that promotes sparsity while robustly handling coefficients of varying magnitudes. A variety of sparsity-promoting priors have been developed, such as the Laplace (Lasso) prior [gelman2013bayesian], the spike-and-slab prior [ishwaran2005spike], or the more flexible Regularized Horseshoe prior [Piironen_2017]. These priors apply shrinkage to irrelevant coefficients while preserving the magnitude of important ones, thereby balancing parsimony and robustness. In particular, the Regularized Horseshoe prior is particularly well-suited for identifying parsimonious dynamical systems, as it combines strong shrinkage for irrelevant coefficients with controlled regularization for significant one; moreover, it avoids the over-shrinkage problem by allowing significant coefficients to escape toward a slab component [hirsh2022sparsifying].
Notably, sparsifying priors alone do not fully guarantee model parsimony. We additionally employ a thresholding step that prunes small-magnitude coefficients after posterior inference. This two-stage approach - probabilistic shrinkage followed by deterministic thresholding - ensures that only the most relevant basis functions are retained, yielding sparse yet reliable governing equations.
3 Methodology
In this work, we introduce BLADE, which incorporates Langevin MCMC methods into the discovery of physical laws in a Bayesian way via the sequential thresholding method. BLADE utilizes the stochastic gradient in the situation of sufficient data collection to improve efficiency and can be exploited for AL in the face of data scarcity (i.e., the time derivative calculation or collection is expensive). We start with the data collected $\mathbf{U}\in\mathbb{R}^{N\times d}$ from the ODE or PDE systems as described before, with the task of identifying the system in terms of (1) and determining the posterior distribution of the model coefficient $\mathbf{\Xi}\in\mathbb{R}^{m\times d}$, where $m$ is the number of candidates in each row of the library $\mathbf{\Theta}(\mathbf{U})$.
3.1 Uncertainty Quantification with Langevin MCMC
For ease of notation, we focus on the column vector $\bm{\xi}\in\mathbb{R}^{m}$ of the model coefficient $\bm{\Xi}\in\mathbb{R}^{m\times d}$, which represents the learned coefficients for one dimension of the system states. This can be easily extended to the sampling of $\bm{\Xi}$ by matching the dimension of the noise term $\bm{\zeta}$ to be introduced.
$d\bm{\xi}_{t}=-\nabla L(\bm{\xi}_{t})\,dt+\sqrt{2\tau}\,d\mathbf{W}_{t},$ (4)
where $L(\cdot)$ represents the energy function, $\{\mathbf{W}_{t}\mid t\geq 0\}$ denotes the standard Brownian motion on $\mathbb{R}^{m}$, and $\tau>0$ is the temperature parameter. In (4), $t$ denotes the continuous algorithmic time of the diffusion process, rather than the physical time of the dynamical system. Under mild conditions on $L$, it is well-established that the diffusion process described by (4) admits a unique strong solution $\{\bm{\xi}_{t},t\geq 0\}$, which is a Markov process. Moreover, as $t\to\infty$, the distribution of $\bm{\xi}_{t}$ converges to the invariant distribution $\pi_{\tau}$, characterized by the density $\pi_{\tau}(\bm{\xi})\propto\exp(-L(\bm{\xi})/\tau)$.
$\tilde{\bm{\xi}}_{k+1}=\tilde{\bm{\xi}}_{k}-\eta_{k}\nabla L(\tilde{\bm{\xi}}_{k})+\sqrt{2\eta_{k}\tau}\,\bm{\zeta}_{k},$ (5)
where $\tilde{\bm{\xi}}_{k}$ indicates a discrete-time approximation of the continuous Langevin diffusion at the $k$-th iteration, $\bm{\zeta}_{k}\sim\mathcal{N}(0,\bm{I}_{m})$ represents independent Gaussian noise, and $\eta_{k}$ is the step size.
To enable scalable and efficient sampling while maintaining convergence to the target distribution under appropriate conditions, we further consider its stochastic gradient version for efficient large-scale data applications. Specifically, consider the following energy function $\tilde{L}$ (negative logarithmic posterior) approximated with a mini-batch of data $\mathbf{B}$ uniformly subsampled from the given data $\mathbf{U}$ [welling2011bayesian]: